What is the approximate peak value of alternating current producing four times the heat produced per second by a steady current of 2A in resistor.???
2.8 A, 4 A, 5.6 A, 8 A

Grade:

2 Answers

Aziz Alam IIT Roorkee
9 years ago
Irms= 2*Isteady(this is given in the question i.e. four times the heat produced)
Irms=Imax/1.414
Using the above 2 relations we can find out that,

Imax=2.828 Isteady
.

Imax=2.828*2=5.6 A
rouf ahmad
9 years ago
i cant understand your 1st equation ,,,can you please eloberate it